OpenStudy (anonymous):

Can you multiply 2x^4 times 6 together to get 12x^4?

OpenStudy (anonymous):

No, the powers would add up each time. 2x^4*2x^4 = 4x^8

OpenStudy (anonymous):

i don't understand your explanation.

OpenStudy (anonymous):

so i can't?

OpenStudy (anonymous):

if 2x = 2*x and 2x^4 = 2*(x^4) => (2*x^4)*6 = 12*x^4

OpenStudy (anonymous):

so i was right? lol

OpenStudy (anonymous):

yeap

OpenStudy (anonymous):

You were right. Sorry I took it that you wanted to multiply the entire number by itself before. I thought it was (2x^4)^6 So, you are CORRECT. :)

OpenStudy (anonymous):

thanks.